# Parity Nonconservation in Rb and Sr+ due to Low-Mass Vector Boson

### v1

@@ Abstract
We calculate the parity non-conserving (PNC) electric-dipole ($E1$) transition amplitudes for the 4s5s4s-5s and $4s-4d_{3/2}$ transitions in Rb and Sr+. Our results include both the nuclear-spin-independent and nuclear-spin-dependent contributions, with particular emphasis on the potential effects of a hypothetical additional $Z^{\prime}$-boson. We highlight possible advantages of using light atoms in searches for such new interaction. The ratio of the contribution of a low mass $Z^{\prime}$-boson to the contribution of the Standard model $Z$ -boson to PNC effects increases rapidly (faster than $1/Z^{2}$) with decreasing nuclear charge $Z$. Another advantage is that theoretical interpretations of experiments in lighter systems may be carried out with a higher accuracy than that in Cs, Ba+, Fr and Ra+.
@@ II PNC amplitudes
where is the Fermi weak interaction constant, $Q_{w}$ is the nuclear weak charge, $\bm{\alpha}$ and $\gamma_{5}$ are the Dirac matrices, $\bm{I}$ is the nuclear spin, and $\rho(\bm{r})$ is the nuclear density normalised to 1. The dimensionless constant $\varkappa$ determines the strength of the SD PNC interaction and is to be found from measurements. The weak nuclear charge $Q_{W}$ is given by [14]
@@
Here $N$ is the number of neutrons, and $Z$ is the number of protons.
@@ III Results and discussion
In both expressions, $N$ is the neutron number, $A$ is the mass number, $A_{\rm PNC}$ is the tabulated numerical value of the spin-independent PNC amplitude expressed in units of $iea_{B}$, see Table 1; $A_{\mathrm{PNC}}^{Z^{\prime}}$ is the tabulated numerical value of the $Z^{\prime}$-boson contribution to the PNC amplitude. The values of $A_{\mathrm{PNC}}^{Z^{\prime}}$ used in Eq. (6) should be taken from the first row of Table 3, while those for Eq. (7) correspond to the last row of the table. Parameter $\varepsilon$ denotes the relative deviation of the experimental PNC amplitude from the Standard Model prediction.

### v2

@@ Abstract
We calculate the parity non-conserving (PNC) electric-dipole ($E1$) transition amplitudes for the 5s6s5s-6s and $5s-4d_{3/2}$ transitions in Rb and Sr+. Our results include both the nuclear-spin-independent and nuclear-spin-dependent contributions, with particular emphasis on the potential effects of a hypothetical additional $Z^{\prime}$-boson. We highlight possible advantages of using light atoms in searches for such new interaction. The ratio of the contribution of a low mass $Z^{\prime}$-boson to the contribution of the Standard model $Z$ -boson to PNC effects increases rapidly (faster than $1/Z^{2}$) with decreasing nuclear charge $Z$. Another advantage is that theoretical interpretations of experiments in lighter systems may be carried out with a higher accuracy than that in Cs, Ba+, Fr and Ra+.
@@ II PNC amplitudes
where $G_{F}$ is the Fermi weak interaction constant, $Q_{W}$ is the nuclear weak charge, $\bm{\alpha}$ and $\gamma_{5}$ are the Dirac matrices, $\bm{I}$ is the nuclear spin, and $\rho(\bm{r})$ is the nuclear density normalised to 1. The dimensionless constant $\varkappa$ determines the strength of the SD PNC interaction and is to be found from measurements. The weak nuclear charge $Q_{W}$ is given by [14]
@@
Here NN is the number of neutrons, and $Z$ is the number of protons.
@@ III Results and discussion
In both expressions, NN is the neutron number, AA is the mass number, $A_{\rm PNC}$ is the tabulated numerical value of the spin-independent PNC amplitude expressed in units of $iea_{B}$, see Table 1; $A_{\mathrm{PNC}}^{Z^{\prime}}$ is the tabulated numerical value of the $Z^{\prime}$-boson contribution to the PNC amplitude. The values of $A_{\mathrm{PNC}}^{Z^{\prime}}$ used in Eq. (6) should be taken from the first row of Table 3, while those for Eq. (7) correspond to the last row of the table. Parameter $\varepsilon$ denotes the relative deviation of the experimental PNC amplitude from the Standard Model prediction.